OpenStudy (anonymous):

What are the next three terms? 80, 20, 5, 5/4, ...?

OpenStudy (anonymous):

Its divided by 4 each time, so what do you think comes next?

OpenStudy (anonymous):

I know this, but I really don't know what comes next

OpenStudy (anonymous):

(5/4)/4 = 5/(4*4) = 5/16

OpenStudy (anonymous):

K. So 5/4 divided by 4 means that the denominator of the fraction is multiplied by 4

OpenStudy (anonymous):

Do do the same for the next two terms

OpenStudy (anonymous):

ok so i have it

OpenStudy (anonymous):

you could also figure out the formula if you want, but you don't have to since it's easy to find the next 3 terms

OpenStudy (anonymous):

5/16, 5/64, 5/256

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

thank you for the help everyone